OpenStudy (anonymous):

A filter circuit is activated when the current reaches 5 amperes. If a charge is fed to the circuit starting at 1 ampere and is steadily increased by 50% of the current value every 10 seconds, how long will it be until the filter activates?

OpenStudy (ineedhelplz):

final value = 5 amperes initial value = 1 ampere rate = 1 + 0.5 = 1.5 every 10 seconds set up a growth equaiton

OpenStudy (ineedhelplz):

\[5=(1.5)^{\frac{ t }{ 10}}\]

OpenStudy (ineedhelplz):

Solve for \[t\]